A bag contains 9 balls numbered 1 through 9. A person draws 4 balls from the bag at the same tome, and then the person writes down the numbers on the balls in increasing order.
What is the probability that the sequence 2, 5, 7, 9 will be written down?
a. 0.00031
b. 0.00794
c. 0.04167
d. 0.22222

Answer by stanbon(75887)

A bag contains 9 balls numbered 1 through 9. A person draws 4 balls from the bag at the same time, and then the person writes down the numbers on the balls in increasing order.
What is the probability that the sequence 2, 5, 7, 9 will be written down?
---
There are 9C4 = (9*8*7*6)/(1*2*3*4) = 126 sequences of 4 digits.
Only 1 is correct.
----
P(correct sequence) = 1/126 = 0.00794
=======
Cheers,
Stan H.
=========

a. 0.00031
b. 0.00794
c. 0.04167
d. 0.22222